OpenStudy (anonymous):

find a quadratic function in standard for that has the following points: (1,-2); (2,-2); (3,-4).

OpenStudy (dumbcow):

plug each point into equation: \[y = ax^{2} +bx +c\] this will give 3 equations with 3 variables (a,b,c) solve system of 3 equations with matrices or elimination

OpenStudy (anonymous):

how do i place them in that equation without an a b and c ?

OpenStudy (dumbcow):

for (1,-2) ..... x=1 and y=-2 ==> -2 = a*(1)^2 + b*1 + c --> -2 = a+b+c

OpenStudy (anonymous):

if i plug it in wouldnt a still be squared or somethin ?

OpenStudy (dumbcow):

a is not squared ... the "x" is squared

OpenStudy (anonymous):

oh. okay so after i do that do i add all of them together ?

OpenStudy (dumbcow):

well you need to know how to solve a system of 3 equations

OpenStudy (anonymous):

what do you mean ?

OpenStudy (dumbcow):

after plugging in the points, you get a + b +c = -2 4a +2b +c = -2 9a + 3b +c = -4 you need to solve for a,b,c

OpenStudy (anonymous):

im trying to write down and figure which numbers can work for all three

OpenStudy (dumbcow):

a=-1 b=3 c=-4

OpenStudy (dumbcow):

y=-1x^2 + 3x -4

OpenStudy (anonymous):

omg thank you. i was going to be thinking forever. and im extra tired. i know i shoulda did it by myself but idk. may you help me with one more question: The weight of a tennis ball must be 57.7 grams, plus or minus 1.7 grams. write an absolute value inequality to illustrate this rule...

OpenStudy (dumbcow):

"Oh my god, thank you. I would have been thinking about this for a long time, and I'm extra tired today. I know that I should have solved it on my own. Can you help me with another problem..." |57.7-x|<1.7

OpenStudy (anonymous):

thank you